Complete and balance the reaction. X Не (aq) Al (OH) (s)3 AICI3(aq)+3H20() HCI(ад) (1 point) Part 2 Water is one of the produ(1 point) Part 3 Write the net ionic equation. (aq), | б. X Не
Answer #1

Ans :

To balance the equation we start by identifying the elements on each side of the equation . Then the number of atoms on one side are equated with thenumber of atoms on the other side.

So the complete and balanced equation is given as :

Al(OH)3 (s) + 3HCl (aq) = AlCl3 (aq) + 3H2O (l)

Part 2 :

Apart from water the other product formed is AlCl3 . The name of this compound is Aluminium chloride.

Part 3 :

The aqueous ions that act as spectator ions in the equation are cancelled out to get the net ionic equation as :

Al(OH)3 (s) + 3H+ (aq) = Al3+ (aq) + 3H2O (l)